# Magnetic effects in the Hadron Resonance Gas

### v1

@@ IV Conclusions
In this work, we examined the effect of finite magnetic field on the thermal properties of hadronic medium. We modeled the hadronic phase of QCD with the hadron resonance gas (HRG) model. For spin-00 particles the densities decrease with magnetic field, while the densities of spin-$1/2$ show the opposite behavior which is due to a different structure of the lowest Landau level. Densities of higher-spin resonances strongly increase with magnetic field which is caused by a singular behavior of the dispersion relation (cf. Eq. (4)) applied to higher-spin particles. Since these densities contribute to final production yields of proton and pions after including the resonance decays, the latter also diverge at sufficiently large fields. This may be problematic for a reliable estimation of magnetic field from non-central HICs based on observed particle yields. As an example, we compared the magnetic field predicted using the HRG and the detailed balance conditions and found a substantial difference between both predictions. Our results suggest that the interpretation of the magnetic field dependence of particle yields in the HRG model, although feasible, should be treated with great care.
We have also studied the second-order fluctuations of conserved charges in conjunction with the recent LQCD data. We find that for our choice of the resonance content of HRG, the LQCD data is underestimated by model results at small fields, while the singular behavior at larger $eB$ leads to overestimation of the data. Although considering hypothetical states improves the description of LQCD data at low $eB$ [31], it makes the breakdown in strong fields even worse.
Therefore, for a reliable comparison between LQCD results and experimental data, it would be crucial to treat the dynamical structure of resonances, especially those with higher spins and neutral states. As the first step, we explored the detailed balance approach and obtained promising results. Another possibility is to generalize the S-matrix approach to finite magnetic field which would allow us to take into account the dynamics of hadrons via the scattering data. Ultimately, non-perturbative approaches capable of describing bound states should be employed to study the dynamics of hadrons in an external magnetic field. Many of these interesting points will be pursued in future research.

### v2

@@ IV Conclusions
In this work, we examined the effect of finite magnetic field on the thermal properties of hadronic medium. We modeled the hadronic phase of QCD with the hadron resonance gas (HRG) model. For spin-00 particles, the densities decrease with magnetic field, while the densities of spin-$1/2$ show the opposite behavior which is due to a different structure of the lowest Landau level. The densities of higher-spin resonances strongly increase with magnetic field which is caused by a singular behavior of the dispersion relation (cf. Eq. (4)) applied to higher-spin particles. Since these densities contribute to final production yields of proton and pions after including the resonance decays, the latter also diverge at sufficiently large fields. This may be problematic for a reliable estimation of magnetic field from non-central HICs based on observed particle yields. As an example, we compared the magnetic field predicted using the HRG and the detailed balance conditions and found a substantial difference between both predictions. Our results suggest that the interpretation of the magnetic field dependence of particle yields in the HRG model, although feasible, should be treated with great care.
We have also studied the second-order fluctuations of conserved charges in conjunction with the recent LQCD data. We find that for our choice of the resonance content of HRG, the LQCD data is underestimated by model results at small fields, while the singular behavior at larger $eB$ leads to overestimation of the data. Although including hypothetical states improves the description of LQCD data at small $eB$ [31], it makes the breakdown in strong fields even worse.
Therefore, for a reliable comparison between LQCD results and experimental data, it would be crucial to treat the dynamical structure of resonances, especially those with higher spins and neutral states. As a first step, we explored the effect of the principle of detailed balance. Another possibility is to generalize the S-matrix approach to finite magnetic field which would allow us to take into account the dynamics of hadrons via the scattering data. Ultimately, non-perturbative approaches capable of describing bound states should be employed to study the dynamics of hadrons in an external magnetic field. Many of these interesting points will be pursued in future research.
Note added. An article discussing similar issues appeared simultaneously with the present work [85].